C= 100 + 0,8Yd

I=150+0,2Y

G=400

T=0,2Y

X=500

M=200+0,24Y

a.      Determining the national equilibrium output and trade balance?

b.     What is the situation of government budget?

c.      If government spending increases 70, how does the output change?

d.     If the exports increase by 100, how does the output change?


1
Expert's answer
2021-03-14T19:09:13-0400

"Y=C+I+G+X-M"

"Y=100+0.8(Y-T)+150+0.2Y+400+500-(200+0.24Y)"

"Y=1150+0.8(Y-0.2Y)+0.2Y-200-0.24Y"

"Y=950+0.64Y+0.2Y-0.24Y"

"Y=950+0.6Y"

"Y-0.6Y=950"

"0.4Y=950"

"Y=2,375" Equilibrium Level of Output

"C=100+0.8(Y-0.2Y)"

"C=100+0.8(0.8*2375)"

"C=100+1,520"

"C=1,620" Equilibrium consumption level

"Imports M= 200+0.24Y \nM= 200+0.24(2375)"

"M=770"

A country's trade balance equals the value of its exports minus its imports.

Trade balance = "X-M"

Trade balance = "500-770"

Trade balance = "-200"

Hence; there is a trade deficit in the country.

b) Situation of Government budget

68% of the budget is funded by the Consumption level, 26% by foreign investments, 19% by government spending.

c) if government spending increases 70, the output will increase by 70 since they are directly related such that an increase in consumption level leads to output by the same amount.

d) If the exports increase by 100 then the output will increase to 2,475 The trade balance deficit will reduce from -200 to -100 making it more favorable.